OpenStudy (anonymous):

Would the degree measure of an angle of 2.25 radians be 128.916° ?

OpenStudy (kfujioka):

Sounds about right.

OpenStudy (mertsj):

Use the conversion factor: 180 degrees = 3.14 radians

OpenStudy (anonymous):

\[\frac{R }{ \pi }=\frac{ D }{ 180^{0} }\] where R is the radians and D is degrees.

OpenStudy (anonymous):

Thank you all!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

your welcome!